differentiate the following function with respect to x
sin(log sin x)

Grade:12

1 Answers

bharat bajaj IIT Delhi
9 years ago
Y = sin( log sin x)
dy/dx = cos(log sinx) d(log sinx)/dx
Dy/dx = cos(log sinx) . Cosx/sinx

Thanks
Bharat Bajaj
Askiitians faculty